What will be the value of Integration of 1/(1-x^2)^2?
 
 

Grade:11

1 Answers

Aditya Gupta
5 years ago
hello ankita. notice that 1/(1-x^2)^2= 1/(1 – x)^2*(1+x)^2
now using partial fractions, write it as
1/(1 – x)^2*(1+x)^2 = A/(1 – x) + B/(1 – x)^2 + C/(1+x) + D/(1+x)^2
now find the constants by ususal methods.
then integrating all these terms separately is easy as they all are standard integrals.